Browsed by
Category: Uncategorized

2 April 2008 – UKIRT and the STFC Programmatic Review – Update 4

2 April 2008 – UKIRT and the STFC Programmatic Review – Update 4

Short-Term Plan

The following note was issued by the Director, JAC, on 2-Apr-2008. 

Dear colleagues, 

I am pleased to announce to the UKIRT community that I have reached agreement with STFC on a new short-term plan for UKIRT. The new arrangements are as follows: 

1. UKIRT will continue operating as a full-service observatory until the end of 2008. This will give me an opportunity to continue my ongoing discussions with several potential partners who are interested in joining UKIRT, some of which are very promising at present. Although some of the candidates are interested only in wide-field imaging with WFCAM, others are interested in the entire suite of capabilities: imaging, spectroscopy and polarimetry. 

2. The situation will be assessed at the end of 2008. Depending on the progress of the partnership discussions and the scientific aspirations of the likely partners, three options are currently foreseeable: 

a) UKIRT will continue operating as a full-service observatory with the UK as a 50% partner; 

b) UKIRT will convert to wide-field survey mode in early 2009 and will continue to operate with the UK as a 50% partner; or 

c) UKIRT will convert to wide-field survey mode in early 2009, the UKIDSS programme will be expeditiously completed, and the facility will then be closed. 

The last option is clearly the worst-case scenario and it is listed here for completeness; I am optimistic that a partnership arrangement can be reached for UKIRT, and that the facility will continue to operate under scenario a or b. 

3. If the situation evolves such that the choice between options becomes clear before the end of 2008, the decision date will be brought forward. 

Please note that the above arrangements assume that UKIRT’s ranking in the Programmatic Review is unchanged; the community submissions to the consultation process are currently being assessed by the ground-based astronomy panel which has been set up for this purpose. The process is described here. If UKIRT’s ranking is significantly altered as a result of the consultation, then the situation would have to be re-asssessed. 

The new plan offers two clear advantages: first, it gives me more time to establish a viable partnership for the future of UKIRT without compromising its capabilities; and second, it recognises the completion of UKIDSS as a minimum deliverable. This is a major step forward for UKIRT and its community. 

Professor Gary Davis
Director, JAC 


Links and History

Previous update 3, 22-Mar-2008STFC Programmatic Review – Update 3: User Input (22-Mar)
Previous update 2, 18-Mar-2008STFC Programmatic Review – Update 2: PPAN Feedback & User Input (18-Mar)
Previous update 1, 10-Mar-2008STFC Programmatic Review – Update 1: Important note to users from Director, JAC and Chair, UKIRT Board (10-Mar)
First note to users, 3-Mar-2008URGENT – STFC Programmatic Review and Semesters 08A and 08B (3-Mar)
25 March 2008 – UKIRT and the STFC Programmatic Review – Update 3

25 March 2008 – UKIRT and the STFC Programmatic Review – Update 3

Please see the latest update, added 02-Apr-2008: STFC Programmatic Review – Update 4: Short-term Plan

Hopefully a final update, added 25-Mar-2008: the mailing list referred to below was closed as planned on the night of 24-Mar. Further emails to this list will receive an automated reply to that effect.


Community Input

The community input form at stfc.ac.uk closed earlier than expected on the 21st. If you had input ready to submit and still wish to have it considered in the Panel’s discussions, please send it by email to stfc_feedback   at   jach.hawaii.edu, from where it will be forwarded to STFC. This address will remain open until 10pm on Monday, 24th March HST. Please try to use text only, and do not include large attachments; and please report any problems to Andy Adamson via the link at the bottom of this page.

22 March 2008 – UKIRT and the STFC Programmatic Review – Updates 2

22 March 2008 – UKIRT and the STFC Programmatic Review – Updates 2

(Update added 22-Mar-2008: the community input period described below is now closed, but see the third update for special circumstances.)


PPAN Project Feedback

Feedback on UKIRT’s priority ranking was received from PPAN today (18-Mar-2007). This underlines the threat to continued operation of UKIRT, and gives the Panel’s rationale for the low science ranking:

UKIRT

We plan to move UKIRT to 100% survey mode as soon as practical after 1 April 2008. Efforts are in hand to find international partners to share the cost of running the telescope in return for access to UKIDSS. Should these fail to materialise, immediate closure would have to be seriously considered.

PPAN recognised that the UKIDSS survey was only partially complete but felt that the overlap with the VISTA survey meant that UKIDSS’ priority was lowered. PPAN also considered that the non-survey science being done at UKIRT could be covered using other telescopes to which UK astronomers had access. Thus, overall, UKIRT fell into the lower priority category.

Category: Lower Priority

User Input to the Consultation Process

In view of the above, we continue to encourage input from UKIRT users to the STFC consultation process (see the link below to go straight to the relevant form, and note that the deadline for submissions is Friday 21-March). Note that comments of the following two types are sought in particular:

  1. Drawing attention to additional important factors that may have been overlooked (in arriving at facility rankings)
  2. Suggesting imaginative ways to maximise the science output in constructing a programme in each subject area, while remaining within a constrained budget.

Our thanks to all those who have already made input. 

Prof. Gary Davis, Director JAC
Dr. Andy Adamson, Associate Director UKIRT


18 March 2008 – UKIRT and the STFC Programmatic Review

18 March 2008 – UKIRT and the STFC Programmatic Review


An important message to UKIRT users: UKIRT and the STFC Programmatic Review

This note contains some essential information for UKIRT operations in semesters 08A and 08B. It also encourages the entire user community, both inside and outside the UK, to respond to the consultation opportunity for the STFC Programmatic Review.

Contents:     1. Background
     2. 08A programmes
     3. 08B Call for Proposals
     4. Response to the consultation exercise
     5. Conclusion

Background

For those not familiar with the funding situation in the UK, here is a brief summary. In 2007, the funding body for astronomy, PPARC, was merged with another research council and part of a third to form a new body, the Science and Technology Facilities Council (STFC). Shortly after its creation, the STFC initiated a Programmatic Review, with the objective of reviewing its entire programme and ranking all commitments in priority order. 

In the autumn, the STFC’s financial allocation from government for the 3-year period starting 1st April 2008 was announced. The net result, once earmarked commitments are subtracted, is a deficit of £80M. Substantial funding cuts are required and there will be considerable pain throughout the programme. The Programmatic Review therefore took on additional significance as the process which would provide the scientific guidance as to where the cuts should fall. 

We are dismayed to have to report that, despite its high productivity and impact and the consequent high profile for UK astronomy, UKIRT was ranked in the lowest of the four priority groups. Ordinarily, this outcome would lead directly to closure of the facility, but STFC have recognised that there are mitigating circumstances which dictate against such precipitate action at this time.

08A Programmes

The consultation document states

“We plan to move UKIRT to 100% survey mode as soon as possible starting 1 April 2008.”

The Board has conferred to consider the implications of this statement, and has adopted the position that, since this is a consultation exercise, it would be inappropriate for the observatory to change its operational plans at the present time. Hence, the current position is that all approved UKIRT 08A programmes will proceed as planned. This includes, inter alia, the Cassegrain block scheduled for July. The situation will be reviewed when STFC’s final programme is developed following the consultation exercise.

08B Call for Proposals

UKIRT’s Call for Proposals for 08B observing programmes was issued on 15th February, and the proposal deadline is 15th March, i.e., within the consultation period. The Board has again decided that it would be inappropriate to change the observatory’s scientific programme in response to a consultation document. Hence, the current position is that the 08B Call for Proposals remains unchanged and we encourage applications for the entire suite of capabilities. The call specified that Cassegrain mode would be offered in August 2008, December 2008 and January 2009. The situation will, again, be reviewed when STFC’s final programme is developed following the consultation exercise. We appreciate that the uncertainty makes it difficult for users to plan their research programmes, but we strongly recommend that you submit your proposals nevertheless. 

Response to the Consultation Exercise

The continued operation of UKIRT is under threat, and we therefore encourage all recent, present and potential users of UKIRT to submit their views. This is an exercise in which weight of numbers may well be important. 

The Board and the UKIDSS consortium are both planning to submit detailed statements to the consultation; the latter will argue for continued operation of the telescope in order to complete the UKIDSS survey programme. In sending this message to the entire user community, we particularly encourage submissions from users who are interested in access to the entire range of UKIRT instrumentation. The proposal demand for the Cassegrain instruments remains healthy, new modes including coronagraphic imaging and polarimetry and full-field thermal infrared imaging are in their infancy, and WFCAM is in demand for non-survey science; it is vitally important that these constituencies make themselves heard. 

The consultation is open to anyone from inside or outside the UK. UKIRT has, since its inception, been open to proposals from any country and the TAG awards time on the basis of scientific merit. UKIRT’s unique capabilities (e.g., in polarimetry) regularly attract a number of international proposals each semester. We therefore strongly encourage submissions to the consultation exercise from non-UK users who value the capabilities offered by UKIRT.

Conclusion

The outcome of the Programmatic Review was surprising and disappointing for UKIRT. The Board and the Observatory intend to respond fully and constructively to the consultation opportunity, and we thank the community of users for assisting us in this. We will endeavour to keep the community informed of developments.

Professor Gary Davis, Director
Professor Pat Roche, Board Chairman
10th March 2008

10 March 2008 – STFC Programmatic Review and UKIRT Semesters 08A & 08B

10 March 2008 – STFC Programmatic Review and UKIRT Semesters 08A & 08B

NOTE: The contents of this page have been superseded by a note from the Director JAC and Chair, UKIRT Board, issued on 10-Mar-2008:

Link


Introduction

STFC has today (3-Mar-2008) issued the outcome of its 2008 Programmatic Review, which now enters a consultation period. The document suggests that STFC plans to switch UKIRT to a “survey-only” mode as soon as possible after April 2008.  The UKIRT Board will meet urgently to discuss this. At present, UKIRT users should assume that the situation for Semesters 08A and 08B remains as follows:

08A

Scheduling of this semester was completed in December and January. Until further notice, users should assume that the Cassegrain block which is scheduled to begin in late June will remain in place. 

08B

The Call for Proposals for 08B, issued on the 15th February, included availability of Cassegrain instruments for August 2008, December 2008 and January 2009. Again until further notice, users should assume that this remains the case. We will of course attempt to clarify this situation ahead of the proposal deadline on 15-Mar. 

20 February 2008 – UKIRT Aluminizing Schedule Altered

20 February 2008 – UKIRT Aluminizing Schedule Altered

Update, 20-Feb-2008

To mitigate the risk of a clash between the late stages of SCUBA-2 commissioning at JCMT and the Aluminizing of the UKIRT primary mirror, the aluminizing has been removed from the telescope schedule for June 2008. It will be rescheduled to fall near the end of August and early September (exact dates to be decided) and the impact on Cassegrain time in August and Wide Field time in the September – November block will be kept to as low a minimum as possible. Additional WFCAM observing nights are being entered into the schedule in June, split in the standard proportions between UKIDSS, Campaign projects, PATT and UH. 

06 February 2008 – Status of UKIRT Due to Poor Weather

06 February 2008 – Status of UKIRT Due to Poor Weather

Update, 6-Feb-2008

UKIRT will remain closed due to bad weather, until 12-Feb-2008. The island of Hawaii has been taking the brunt of a series of winter weather systems and the summit has been continually snowed in for many days. The length of the closure is determined by the need to carry out Cassegrain-to-WFCAM changeover engineering over a protracted period of short working days.

02 January 2008 – WFCAM Stop Press

02 January 2008 – WFCAM Stop Press

NOTE – information on accessing WFCAM data sets is now available via the normal UKIRT data access page.


Jan 2008

Final update: this page is obsolete and will no longer be used to record WFCAM status.     – AJA


27-Mar-2006

Update: WFCAM has operated smoothly in recent weeks. Improvements in focus tracking implemented by Tom Kerr have improved image quality stability and since new year the intrinsic seeing on the mountain has improved such that we have been able to take advantage of these improvements.


01-Nov-2006

WFCAM is back on the telescope, has checked out ok and survey operations have begun again. Efficiency savings resulting from the acquisition system upgrade carried out over the summer appear to be significant; 10 second exposures are now of order 80% efficient. More improvements are possible and this page will be updated as things develop.


21-Apr-2006

WFCAM is cold and all four arrays have checked out ok today.


18-Apr-2006

Preparation for the upcoming WFCAM block is in full swing. Database modifications have been made which will allow moon avoidance for the UKIDSS surveys, and we are in the process of building up new, faster PCs for the acquisition system. Top to bottom tests of one of these systems, from the OMP through to data on disk, should be possible this week.


1-Mar-2006

WFCAM remains off the telescope. Work is ongoing at the UK ATC and in Hilo to upgrade the acquisition systems (goal to improve acquisition overheads).


10-Feb-2006

WFCAM has just come off the telescope after completion of a successful 5 month observing block. It goes back on at the end of April. The forward-look schedule details expected observing blocks through 2007.


4-Nov-2005

WFCAM remains in smooth operation, hence the lack of updates since the end of August. There has been one issue with image elongation in the last week or so, the root cause of which appears to be astigmatism in the primary mirror; wavefront sensing data taken last night should be sufficiently good to determine the cause of this and remove it.


31-Aug-2005

WFCAM is working well. Wavefront sensing results from the engineering nights are being analysed. Telescope oscillations have been addressed  recently; the fast oscillation which afflicted observations in the North have been essentially removed via  some adjustments to the mount servo parameters (additional work on the mount electronics will be done when time permits). The large, slow oscillation which was largely compensated for by the secondary (but which caused occasional losses of guiding when doing large offsets) has been traced to a software problem in the link between the MPIA crate and mount computer; this was  fixed this evening and the  oscillation appears to  be gone.  The optimistic assessmnt of automatic guide star acquisition which resulted from the first engineering night has been confirmed.


26-Aug-2005

Notes from a more extended first night than expected (JAC ETS got WFCAM on the telescope and cold very quickly and we took advantage of this to get an early start on both the guider and the infrared arrays):

  • Russell’s Automatic guide star aquisition works well.
  • Initial pointing model measured and implemented. Pointing will change as mirror parameters are tuned, so will do a full pointing test following optical tuning.
  • Three focal plane tilt measurements, with M2 at the boresight (U=2.8, V=3.2), indicated (very small) average tilts of: EW: 9.0 microradians NS: -197.8 microradians.
  • NS tilt was corrected by tilting M2 and the residual came down to the <50 microradian level at zenith.
  • In this configuration, image quality of 0.65arcsec was delivered at H on all 4 arrays. Gemini reporting 0.5″ in V at that point. Looking at out of focus images, some astigmatism and other aberrations are apparent. A wavefront sensing run will be done tomorrow night.
  • Made predicted adjustments to the focus offsets to compensate for the AG filters that were swapped out during the engineering, and did some coarse tuning. Focus offsets for both the IR and AG side for all filters are now pretty close, but will need further optimisation over the next few nights. This will become easier as the instrument cryogenic temperatures stabilise.
  • Did a handful of standard fields as they should be useful (it’s a photometric night apparently) and they’re good excercise of many parts of the system. Automatic guide star aquisition handled them nicely – it was possible to queue up about 5 standard fields and the system ran through them all, doing 5 filters on each, without any intervention from either observer or TSS, apart from a pointing tweak right at the start. This bodes well for survey efficiency.
  • A standard field was observed with the field lens exposed to the (~50% illuminated) moon. The large patches on the lens have been sucesfully removed, though there are a few dust specks or similar, looks worst in the NW corner but probably acceptable.

23-Aug-2005

WFCAM is being installed on the telescope this week. The first observing night (for guider and other checkout in advance of the infrared arrays being fully cold) will be Thursday 25th.


16-Aug-2005

Cold tests were carried out yesterday – WFCAM is cold, and the mechanisms, all 4 IR arrays and AG CCD are all present and correct. Autoguider noise appears slightly elevated but this is on the dome floor and is known to behave differently when on the telescope.


1-Aug-2005

The 05B UKIRTOT and UKIRTSDT will be released on the 2nd August HST.


28-Jul-2005

WFCAM internal tilt shims were adjusted yesterday. The instrument is now closed up again and pumping down. Warm mechanism tests were carried out successfully. Cooldown is scheduled for the week after next.


11-Jul-2005

Preparatory calculations for the internal tilt adjustment will be under way shortly, employing all relevant data obtained in the course of the previous observing block. Field lens cleaning (rear surface) is scheduled for late July. This replaces the scheduled installation of the flat field screen, which has been delayed prior to shipping. The screen is still expected to be installed before WFCAM goes back on the sky in late August.


24-Jun-2005

WFCAM is now off the telescope until late August. Work to be done while it is off includes:

(i)  cleaning of the inaccessible rear surface of the field lens
(ii)  further adjustment of internal tilt at the focal plane
(iii)  testing of data reduction machine RAID configurations
(iv)  assessment and possible fixing of per-frame observing overheads

Work continues on DR pipeline. The note of 2-Jun  was overoptimistic. The summit pipeline was not capable of keeping up with data taking in all modes used by UKIDSS. 


2-Jun-2005

Recent data indicate that following from adjustment of spherical aberration undertaken on the basis of wavefront sensing measurements reported in the previous posting, the image quality floor has moved to smaller FWHM, with 0.6 arcsecond images now occurring with reasonable regularity. More work remains to be done, including a further internal adjustment when the cryostat comes off the telescope later this month.

Considerable improvements have been made to the CASU summit pipeline and we are able to reliably monitor sky brightness and image quality through the night, in real time.


18-May-2005

Cryostat tilt was altered on Monday this week, using the accumulated information on tilt from engineering over the previous week. The required secondary mirror tilt to remove the remaining focal plane tilt was seen to decrease accordingly. A wavefront sensing scheduled for last night was cancelled due to winds in excess of the level normally usable for these measurements; this will be repeated tonight.


9-May-2005

Image quality currently floors at 0.7 arcseconds, and efforts to understand and remove the cause continue. Wavefront sensing results are better understood following from systematic WFS measurements taken on the 4th of May. The night of 11-May will be put into further sustematic tests of image quality and focal plane tilt as a function of M2 decentre and tilt. The night of the 16th will be devoted to image quality tests with the whole cryostat tilt having been adjusted externally. 


2-May-2005

Additional wavefront sensing observations will be taken tonight, to provide further data on the effect of secondary mirror positioning on the delivered image quality.  The results will be analysed in conjunction with the UK ATC.


26-Apr-2005

Update: the recent scheduled service half nights were put toward further image quality engineering observations. Service projects will now be undertaken on the nights of the 4th and 5th of May, flexed with engineering and UKIDSS.


19-Apr-2005

Apologies for the gap due to AJA’s absence from the island. Summary of current status:

After losing the first four nights to weather, a compressed schedule was adopted including two nights testing the effect of tilting the cryostat relative to the telescope. The results were a convincing indication that the default position was best, though we have not been able to test this exhaustively (essentially measured image quality at two extreme east-west positions and found them worse than at the default in one or more respects). 

Calibration and SV observations have now been carried out for a number of nights, interspersed with scheduled UH runs. The system is working quite well, with two main outstanding issues: (i) an acquisition overhead of a few seconds per frame which is being investigated (it is probably a software issue) and (ii) an apparent 0.1-0.2 arcsecond final delivered image quality deficit which will be subjected to further wavefront sensing analysis in engineering opportunities over the coming week.


1-Apr-2005

First two commissioning nights have been significantly affected by thick clouds but it has been possible to find a combination of secondary tilt which removes the residual focal plane tilt. Conditions have not yet been good enough to check for K background levels with this alignment. Wavefront sensing measurements have been made and the resulting adjustment of the primary has improved image quality. All the preceding notes should be  interpreted in the light of the rather thick cloud cover through which most of the measurement s have been taken.


30-Mar-2005

WFCAM is in place on the telescope and cooling back down. First night of commissioning observing will be Thursday (31-Mar), though weather is looking less than ausipcious for the first two nights. 

WFCAM has been added to the service system database and we are now formally accepting WFCAM service proposals. More details in the news item on the home page news section.


29-Mar-2005

Cooldown has been in progress since Monday, with no reported problems to date. Expect WFCAM on the telescope on Wednesday, balance and ready by Thursday evening as scheduled. Field lens inspection was done last week and all appeared  clean. Will be reinspected this week before putting  in place on the cryostat.



23-Mar-2005

Rework on the focal plane tilt was completed yesterday and the cryostat is now pumping down in preparation for cooldown next week. Schedule remains the same: back on sky on the 31st March (for initial engineering checkout; internal components will still be contracting at that point, expect first useful information on focus etc. to come on the 1st April.


21-Mar-2005

Additional information from wavefront sensing and background measurements taken on the engineering night on the 16th: once the focal plane tilt is finally removed, the secondary tilt position which produces best image quality will also be that which minimizes the K band background. Schedule remains as given on the 14th March.


17-Mar-2005

The night of 16-Mar HST was usable (cirrus, moon) and various test engineering observations were taken. The autoguider appears to be performing very well; guiding at 30Hz was possible on a 17th-magnitude star in thin cirrus and with the moon up. An sdt test at 12-13 hours, +30 degrees shows that with a 17-mag limit, backtracking is necessary on 20 cases out of 160 pointings (and the whole area is successfully tiled); with 17.5-mag, which seems reasonable given the poor conditions of the guiding test, backtracking is required only 9 times. Summary: tiling the galactic cap will be possible.


14-Mar-2005

The East-West tilt apparent in the first night of WFCAM data (since which we have been closed due to weather) is confirmed, suggesting that some fraction of the shimming in that direction was applied incorrectly. Since all but one of the six scheduled engineering nights have now been lost to weather, the time available for calibration setup and science verification has now been reduced to uselessly small levels, and the likelihood of continued poor weather makes the plan outlined in the note of 16-Feb-2005 (below) ineffective and probably wasteful of time in April. That plan has been abandoned in favour of an alternative in which the cryostat engineering is done while the weather remains poor.

The cryostat will be removed on Thursday 17-Mar and warmed up for rearrangement of the shimming. Given the current weather, Wednesday night (16-Mar) is the only night on which any observing may be possible; this will be devoted to completion of work on the Autoguider, to taking flat fields with the new infrared readout, and to some exploratory wavefront-sensing observations. The anticipated schedule is approximately as follows:

17-Mar Cryostat off and warmup initiated
21-Mar Open-cryostat engineering (week)
23-Mar Pump
28-Mar Cooldown initiated
31-Mar First on-sky observing (still cooling)
1-Apr – 7-Apr Engineering observations to completion; calibration observing commences
8-Apr – 12-Apr [First halves – UH run] : [Second halves – UKIDSS SV and calibration observations]

After this point the schedule remains as originally posted, although some rearrangement of run blocks may be made with a view to generating more contiguous blocks of WFCAM time for UKIDSS startup and calibration verification. Any affected observers will be contacted directly.

Note: to ensure a quick turnaround, we will not be reverting to Cassegrain observing while WFCAM is being worked on.


13-Mar-2005

A possible problem has been recognized with the focal plane tilt measurement referred to on 11-Mar. It is still to be confirmed but we cannot at this point rule out the posibility that a further adjustment to the tilt needs to be made. We will issue an update on Monday.


11-Mar-2005

First on-sky night proved much better than anticipated on the basis of the weather forecast; a large dataset was amassed on the focal plane alignment and guider focus offsets. Initial indications are that the focal plane is now perpendicular to the optical axis to as well as it is possible to measure. Work is continuing on the guider offsets and final statements will be posted on this page. Data are currently being taken with the secondary not powered up, and statements are subject to change.


9-Mar-2005

Cryostat, field tower and secondary are safely installed on the telescope and the cryostat is cooling back down. Tomorrow night (10-Mar-2005 HST) will, as planned, be the first night of on-sky commissioning. Weather conditions do not look positive for the first few nights at least; a front is due to stall over the Big Island and various local precipitation is anticipated. If this proves to be the case, and assuming that no major optical problem is discovered in the early stages, the priority list of (1) basic engineering observing (focal plane tilt, focus offsets, wavefront sensing, optical alignment etc.), (2) Calibration fields, and (3) UKIDSS science verification will be redistributed, to cover a period up to and including 28-Mar-2005.  

Derek Ives and Paul Hirst have been working on readout characteristics of the infrared arrays; the reset anomaly has been reduced by a factor of approximately 10 and appears to be more stable. More information on performance will be posted once the instrument sees sky.


7-Mar-2005

The instrument will be installed on the telescope over the coming three days. First on-sky commissioning night will be 10-Mar-2005. All appears well with the hardware and significant improvements have been made on the infrared array readout, reset anomaly and the autoguider sensitivity in the past few days. An update will be issued on this page when  details have been verified quantitatively on sky.


28-Feb-2005

WFCAM appears to be in good shape – cold tests done this morning show the mechanisms and all four arrays functional. It will go on the telescope as scheduled on the 7th.


23-Feb-2005

WFCAM Cooldown continues.
Survey Definition Tool released.


22-Feb-2005

WFCAM Cooldown started on schedule this morning.


18-Feb-2005

Schedule update: cold test information should be comprehensive enough by 28th Feb to give a “go/no go” decision on installing on the telescope. UKIDSS will be informed on that day (HST). All other schedule details remain as stated on the 16th.

SDT: the Survey Definition Tool requires some repackaging before final release. This is being worked on between JAC and UKATC. the Observing Tool is already released (see the Prep/Exec page for a link to the download page).


16-Feb-2005

Update on schedule following JAC-ATC telecon today:

WFCAM installation on telescope – 7-9 March
WFCAM initial on-sky engineering – 10-22 March

Option 1 – no problems uncovered in engineering time

WFCAM engineering, queue and service nights continue through to 28-Mar, with PATT, UKIDSS and UH runs commencing on the 29th.

Option 2 – optical problems discovered in on-sky engineering time are sufficient to require focal plane adjustment

Whether this option is needed will be clear by 16-Mar, from measurements undertaken in the first week. In this case, the cryostat will be warmed up on the 23rd and back on sky on the 8th April; there will follow a checkout and alignment period of approximately one week, before the commencement of normal observing programmes on approximately 13-Apr. 


15-Feb-200523-Mar-2005

Work on the focal plane tilt was completed yesterday and the cryostat is now pumping down in preparation for cooldown next week. Schedule remains the same: back on sky on the 31st March (for initial engineering checkout; internal components will still be contracting at that point, expect first useful information on focus etc. to come on the 1st April.


21-Mar-2005

Additional information from wavefront sensing and background measurements taken on the engineering night on the 16th: once the focal plane tilt is finally removed, the secondary tilt position which produces best image quality will also be that which minimizes the K band background. Schedule remains as given on the 14th March.


17-Mar-2005

The night of 16-Mar HST was usable (cirrus, moon) and various test engineering observations were taken. The autoguider appears to be performing very well; guiding at 30Hz was possible on a 17th-magnitude star in thin cirrus and with the moon up. An sdt test at 12-13 hours, +30 degrees shows that with a 17-mag limit, backtracking is necessary on 20 cases out of 160 pointings (and the whole area is successfully tiled); with 17.5-mag, which seems reasonable given the poor conditions of the guiding test, backtracking is required only 9 times. Summary: tiling the galactic cap will be possible.


14-Mar-2005

The East-West tilt apparent in the first night of WFCAM data (since which we have been closed due to weather) is confirmed, suggesting that some fraction of the shimming in that direction was applied incorrectly. Since all but one of the six scheduled engineering nights have now been lost to weather, the time available for calibration setup and science verification has now been reduced to uselessly small levels, and the likelihood of continued poor weather makes the plan outlined in the note of 16-Feb-2005 (below) ineffective and probably wasteful of time in April. That plan has been abandoned in favour of an alternative in which the cryostat engineering is done while the weather remains poor.

The cryostat will be removed on Thursday 17-Mar and warmed up for rearrangement of the shimming. Given the current weather, Wednesday night (16-Mar) is the only night on which any observing may be possible; this will be devoted to completion of work on the Autoguider, to taking flat fields with the new infrared readout, and to some exploratory wavefront-sensing observations. The anticipated schedule is approximately as follows:

17-Mar Cryostat off and warmup initiated
21-Mar Open-cryostat engineering (week)
23-Mar Pump
28-Mar Cooldown initiated
31-Mar First on-sky observing (still cooling)
1-Apr – 7-Apr Engineering observations to completion; calibration observing commences
8-Apr – 12-Apr [First halves – UH run] : [Second halves – UKIDSS SV and calibration observations]

After this point the schedule remains as originally posted, although some rearrangement of run blocks may be made with a view to generating more contiguous blocks of WFCAM time for UKIDSS startup and calibration verification. Any affected observers will be contacted directly.

Note: to ensure a quick turnaround, we will not be reverting to Cassegrain observing while WFCAM is being worked on.


13-Mar-2005

A possible problem has been recognized with the focal plane tilt measurement referred to on 11-Mar. It is still to be confirmed but we cannot at this point rule out the posibility that a further adjustment to the tilt needs to be made. We will issue an update on Monday.


11-Mar-2005

First on-sky night proved much better than anticipated on the basis of the weather forecast; a large dataset was amassed on the focal plane alignment and guider focus offsets. Initial indications are that the focal plane is now perpendicular to the optical axis to as well as it is possible to measure. Work is continuing on the guider offsets and final statements will be posted on this page. Data are currently being taken with the secondary not powered up, and statements are subject to change.


9-Mar-2005

Cryostat, field tower and secondary are safely installed on the telescope and the cryostat is cooling back down. Tomorrow night (10-Mar-2005 HST) will, as planned, be the first night of on-sky commissioning. Weather conditions do not look positive for the first few nights at least; a front is due to stall over the Big Island and various local precipitation is anticipated. If this proves to be the case, and assuming that no major optical problem is discovered in the early stages, the priority list of (1) basic engineering observing (focal plane tilt, focus offsets, wavefront sensing, optical alignment etc.), (2) Calibration fields, and (3) UKIDSS science verification will be redistributed, to cover a period up to and including 28-Mar-2005.  

Derek Ives and Paul Hirst have been working on readout characteristics of the infrared arrays; the reset anomaly has been reduced by a factor of approximately 10 and appears to be more stable. More information on performance will be posted once the instrument sees sky.


7-Mar-2005

The instrument will be installed on the telescope over the coming three days. First on-sky commissioning night will be 10-Mar-2005. All appears well with the hardware and significant improvements have been made on the infrared array readout, reset anomaly and the autoguider sensitivity in the past few days. An update will be issued on this page when  details have been verified quantitatively on sky.


28-Feb-2005

WFCAM appears to be in good shape – cold tests done this morning show the mechanisms and all four arrays functional. It will go on the telescope as scheduled on the 7th.


23-Feb-2005

WFCAM Cooldown continues.
Survey Definition Tool released.


22-Feb-2005

WFCAM Cooldown started on schedule this morning.


18-Feb-2005

Schedule update: cold test information should be comprehensive enough by 28th Feb to give a “go/no go” decision on installing on the telescope. UKIDSS will be informed on that day (HST). All other schedule details remain as stated on the 16th.

SDT: the Survey Definition Tool requires some repackaging before final release. This is being worked on between JAC and UKATC. the Observing Tool is already released (see the Prep/Exec page for a link to the download page).


16-Feb-2005

Update on schedule following JAC-ATC telecon today:

WFCAM installation on telescope – 7-9 March
WFCAM initial on-sky engineering – 10-22 March

Option 1 – no problems uncovered in engineering time

WFCAM engineering, queue and service nights continue through to 28-Mar, with PATT, UKIDSS and UH runs commencing on the 29th.

Option 2 – optical problems discovered in on-sky engineering time are sufficient to require focal plane adjustment

Whether this option is needed will be clear by 16-Mar, from measurements undertaken in the first week. In this case, the cryostat will be warmed up on the 23rd and back on sky on the 8th April; there will follow a checkout and alignment period of approximately one week, before the commencement of normal observing programmes on approximately 13-Apr. 


15-Feb-2005

WFCAM has been reassembled, closed up and is pumping down in preparation for cooldown which will be done during week of 21-Feb. Current schedule suggests a one-week slip can be held to, with the camera going back on the telescope on the 7th-9th March and on-sky commissioning commencing on 10-Mar.


8-Feb-2005

Good news on the array 4 problem (based on report by E.Starman).  The detector flexi cable was changed today and the channel 9 problem is fixed.  All 4 detectors and CCD checked ok with the spider assembly out of the instrument.  Recheck of filter paddles/detector box fit and instrument reassembly scheduled for tomorrow and Thursday. Current expectation is that the cooldown will commence on 21-Feb-2005.


5-Feb-2005

WFCAM detector 4 was found to have a bad channel on closing up after focal plane tilt adjustment.

Engineering to fix this is currently scheduled for Tuesday 8-Feb-2005. We expect at least one weeks’ delay in WFCAM being on sky (shutdown for installation now scheduled for 7-Mar; first commissioning night 10-Mar). The schedule on the web will be updated once observers have been allocated for the gap which has opened up.

WFCAM has been reassembled, closed up and is pumping down in preparation for cooldown which will be done during week of 21-Feb. Current schedule suggests a one-week slip can be held to, with the camera going back on the telescope on the 7th-9th March and on-sky commissioning commencing on 10-Mar.


8-Feb-2005

Good news on the array 4 problem (based on report by E.Starman).  The detector flexi cable was changed today and the channel 9 problem is fixed.  All 4 detectors and CCD checked ok with the spider assembly out of the instrument.  Recheck of filter paddles/detector box fit and instrument reassembly scheduled for tomorrow and Thursday. Current expectation is that the cooldown will commence on 21-Feb-2005.


5-Feb-2005

WFCAM detector 4 was found to have a bad channel on closing up after focal plane tilt adjustment.

Engineering to fix this is currently scheduled for Tuesday 8-Feb-2005. We expect at least one weeks’ delay in WFCAM being on sky (shutdown for installation now scheduled for 7-Mar; first commissioning night 10-Mar). The schedule on the web will be updated once observers have been allocated for the gap which has opened up.

16 November 2007 – Status of UKIRT Top-End Fault

16 November 2007 – Status of UKIRT Top-End Fault

Update 16-Nov-2007

UKIRT resumed normal operation tonight; the fault appears to be essentially cured and image quality is looking nominal. We hope that this will be the final update on this topic.

Update 14-Nov-2007

A tip-tilt piezo actuator failed due to water damage occurring during bad weather in early November. This fault has proved multifaceted, difficult to remedy and work is still in progress. This page will be updated as more news becomes available.